OpenStudy (anonymous):

MEDAL!!! The leg of a right triangle is 3 units and the hypotenuse is 11 units. What is the length, in units, of the other leg of the triangle?

OpenStudy (anonymous):

@AndrewTheCookie

OpenStudy (turingtest):

do you know the pythagorean thoerem?

OpenStudy (anonymous):

i dont understand it though. I'm not good at math at all!

OpenStudy (turingtest):

the pythagorean theorem says that[\c^2=a^2+b^2\]which side is the hypotenuse?

OpenStudy (turingtest):

\[c^2=a^2+b^2\]

OpenStudy (anonymous):

the right side.

OpenStudy (anonymous):

square root 112?

OpenStudy (turingtest):

no i meant in the picture i drew, which letter is the hypotenuse?

OpenStudy (anonymous):

@TuringTest

OpenStudy (anonymous):

a

OpenStudy (turingtest):

no, it is c the hypotenuse is the longest side of a right triangle

OpenStudy (turingtest):

so put the numbers given in the problem into the formula\[c^2=a^2+b^2\]

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

11^2-3^2

OpenStudy (alexandervonhumboldt2):

by pyth theorem

OpenStudy (anonymous):

132 @mathstudent55

OpenStudy (alexandervonhumboldt2):

121-9=112

OpenStudy (alexandervonhumboldt2):

112

OpenStudy (alexandervonhumboldt2):

@chosenmatt

OpenStudy (anonymous):

square root?